OpenStudy (anonymous):

Calculate the fractional average of the numbers 1/2, 5/8 and 7/16

OpenStudy (anonymous):

add them then divide by 3

OpenStudy (anonymous):

how you mean? Can you show me?

OpenStudy (phi):

can you find the common denominator of these fractions?

OpenStudy (anonymous):

no

OpenStudy (phi):

the common denominator is 16. if you knew 1/2 is the same as 8/16 (it is) and you add 8/16 + 7/16 ?

OpenStudy (anonymous):

15/16

OpenStudy (phi):

now add 5/8 to that. First change 5/8 to 10/16 add to 15/16

OpenStudy (anonymous):

25/16

OpenStudy (phi):

now divide 25/16 by 3 \[ \frac{25}{16} \div 3 \] the way to divide is "invert" the 3 and multiply. That means flip 3/1 to make it 1/3 and multiply: \[ \frac{25}{16} \div 3= \frac{25}{16}\cdot\frac{1}{3} \] multiply top times top and bottom times bottom